OpenStudy (anonymous):

Consider the line that passes through the point (-1, 6) and has a slope of -4. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

jigglypuff314 (jigglypuff314):

point-slope form is y-y1=m(x-x1) slope-intercept form is y=mx+b standard form is ax+by=c

jigglypuff314 (jigglypuff314):

then you just plug in the values Part 1: y-y1=m(x-x1) Given: (-1,6) => (x1,y1) so x1=-1 and y1=6 slope = -4 = m Part 2: using what you get from Part 1 you can distribute and add 6 to both sides Part 3: you can add 4x to both sides to what you got in Part 2